Please expert solve this quick ?

TheWheat producers of North America are expecting to match their supply to meet the change in the market demand during coming year that is expected to change. The current year market supply and supply function is assumed to be:

QD= 62.5 - 0.125P

QS= 0.5P - 100

Please predict either a rise or fall in the demand of wheat as a percentage change from the current year demand based on your observations from the website and derive a new demand function with the predicted percentage change (EITHER INCREASE OR DECREASE) in demand, (i.e. the quantity demanded will increase or decrease by THE PERCENTAGE for each level of price). (Please restrict your change in demand within ± 25%)

Please compute the following:

  1. Market Equilibrium Price for current year and next year with percentage change.
  2. Market Equilibrium Quantity for current year and next year with percentage change
  3. Producer Surplus and Consumer Surplus for current year and following year with percentage change.

Please keep this info in mind:

the function then decreases by 20% leading for the new function to become Qd = 50 - 0.1P
Answer #1

soal D n 259 c g Demond 25 30 30 62.5 quercity

Qd = 62.5 - 0.125P

Qs = 0.5P - 100

At equilibrium, demand = supply

62.5 - 0.125P = 0.5P - 100

162.5 = 0.625P

P = 260

At this P, Q = 30

If demand curve changes to Qd = 50 - 0.1P

New equilibrium: 50 - 0.1P = 0.5P - 100

150 = 0.6P

P = 250

At this P, Q = 25

New equilibrium point occurs when D1 = supply. At this point, price is 250 and quantity traded is 25 units.

a) Price of this year = 260

Price of next year = 250

%change in price = [(250 - 260) / 260] * 100 = -3.84%

b) Quantity of current year = 30

Quantity of next year = 25

%change in Quantity = [(25 - 30 / 30] * 100 = 16.67%

c) Consumer surplus in current year is area of portion A + D whose sum is (1/2) * (500 - 260) * (30 - 0) = 3,600

Producer surplus in current year is area of portion B + C + E whose sum is (1/2) * (260 - 200) * (30 - 0) = 900

Consumer surplus in next year is area of portion A + B whose sum is (1/2) * (500 - 250) * (25 - 0) = 3,125

Producer surplus in next year is area of portion C whose sum is (1/2) * (250 - 200) * (25 - 0) = 625

%change in consumer surplus = [(3,125 - 3,600) / 3,600] * 100 = -13.19%

%change in producer surplus = [(625 - 900) / 900] * 100 = -30.55%